# Nonlinear trident using WKB and worldline instantons

### v1

@@ Abstract
We consider nonlinear trident, $e^{\scriptscriptstyle-}\to e^{\scriptscriptstyle-}e^{\scriptscriptstyle-}e^{\scriptscriptstyle+}$, in various electric background fields. This process has so far been studied for plane-wave backgrounds, using Volkov solutions. Here we first use WKB for trident in time-dependent electric fields, and then for fields which vary slowly in space. Then we show how to use worldline instantons for more general fields which depend on both time and space.
@@ I Introduction
There continues to be growing progress in strong-field-QED processes, especially those occurring when electrons or photons collide with high-intensity laser fields [1, 2, 3]. One process is the nonlinear trident process, $e^{\scriptscriptstyle-}\to e^{\scriptscriptstyle-}e^{\scriptscriptstyle-}e^{\scriptscriptstyle+}$. Many particles can be produced by a single electron, because factors of $\alpha=1/137$, which would otherwise suppress higher-order processes, are compensated for by large factors due to the high intensity or extended pulse length of the background field. To theoretically model such higher-order processes, one needs to employ various approximations. One approximation, in particular, is to treat higher-order processes as incoherent product of first-order processes. One reason that we are interested in studying trident is that it is the first nontrivial step in the production of more particles while still being sufficiently “simple” so that one can calculate and check the corrections to the incoherent-product approximation. The calculations of these corrections are actually quite nontrivial. In fact, while the simplest part of the correction (the direct part) was calculated already in the 70’s by [4, 5], it was not until [6, 7] that the rest of the correction (the exchange part) was calculated. It turned out that this previously omitted exchange term is in fact on the same order of magnitude as the simpler direct term. Here we should mention that there have also been fully numerical computations of the trident spectrum which did not worry about separating the probability into an incoherent product and corrections to it. A not complete list of papers studying trident is [4, 5, 8, 9, 10, 11, 12, 6, 7, 13, 14, 16, 15, 17, 18].
All these previous studies of trident have in common that they considered the background field to be a plane wave, i.e. with a gauge field given by [^1] We absorb the charge into the definition of the background field, $eF^{\mu\nu}\to F^{\mu\nu}$, and set $c=\hbar=m=1$, where $m$ is the electron mass. $A_{\mu}(kx)$, where $k_{\mu}=\omega(1,0,0,1)$ is the wave vector and $kA=0$. There are good reasons for this. One is the fact that an electron with sufficiently high energy tends to see a quite general background as if it were a plane wave. However, if the electron’s momentum is (approximately) parallel an electric field, then the electron will not see that field as a plane wave. And if the field is sufficiently strong to significantly bend the electron trajectory, then one can also expect differences from the plane-wave approximation. Then there is of course cases where the electron energy is not high.
@@
In this paper we will use weak-field approximations to calculate the trident probability for non-plane-wave backgrounds. We will start with a WKB treatment of trident in purely time-dependent electric fields, $E(t)=A_{3}^{\prime}(t)$. One might argue about whether this can be seen as going “beyond” plane waves. We are not saying that fields on the form $E(t)$ actually give a more accurate model of realistic laser fields. For one thing, $E(t)$ does not solve Maxwell’s equation, while plane waves do. However, from an analytical point of view, there is a sense in which considering $E(t)$ is going beyond plane waves. The fact that the two field invariants, ${\bf E}^{2}-{\bf B}^{2}$ and ${\bf E}\cdot{\bf B}$, are both zero for plane waves, means that the probabilities have less structure. For $E(t)$, on the other hand, one of the field invariants is nonzero, so even before doing any calculations, one can expect the result to have more structure. In fact, results for a general plane wave can be obtained from the results for $E(t)$ by taking the limit where the transverse momentum of the electron is very high. In other words, the plane-wave results are contained in the results for $E(t)$, so in this sense we are going beyond plane waves when studying $E(t)$. Considering $E(t)$ therefore allows us to get a sense of what is neglected when approximating the background as a plane wave, and when it is not possible to do so. We show, in particular, that if we instead take the limit where the momentum of the electron is large but parallel to the electric field, then one does not recover the plane-wave result.
@@
Then we show how to use worldline instantons to study trident in general space-time dependent fields. Both the WKB and the instanton approach calculates the same, leading order in a weak-field approximation. In the WKB approach one first have to solve the Dirac equation, i.e. one has to find an approximation of a field, $\psi(x^{\mu})$, which in general depends nontrivially on $t$, $x$, $y$ and $z$ in addition to the field and momentum parameters, and then one has to plug that solution into the Feynman rules and perform the integrals (which one can do with the saddle-point approximation). Note that $\psi(x^{\mu})$ contains more information than what is actually needed for any particular process. In contrast, in the instanton approach one goes directly for the amplitude of the particular process one considers, i.e. without first calculating some intermediate result. Moreover, the instantons are 1D trajectories which permits a more intuitive understanding of these processes. With this in mind, it is perhaps not so surprising that we can calculate things with the instanton approach which, although in principle possible, would be difficult to do with a WKB approach. We will show how to apply the instanton approach for fields which depend on both time and space separately (i.e. not just on e.g. $t-z$ as in the plane-wave case). This can be seen as going beyond plane waves both in the analytical sense described above as well as in a more direct sense of more accurately modeling realistic lasers.
@@
  • We want develop methods to be able to calculate the probability of trident and similar processes for fields which depend on more than one space-time coordinate. Our open-worldline approach seems promising for that.
@@ II Trident in $E(t)$ using WKB
We first consider time-dependent fields on the form $A_{3}(-t)=-A_{3}(t)$, $A_{3}(t)=f(\omega t)/\gamma$, where $f$ is a dimensionless function and $\gamma=\omega/E$ is the Keldysh parameter. In the high-energy limit we will often express things in terms of $a_{0}=1/\gamma$ (the classical nonlinearity parameter) instead of $\gamma$, because $a_{0}$ is more common when dealing with plane-wave backgrounds. We use the following notation for the momentum, $\pi_{\scriptscriptstyle\perp}=q_{\scriptscriptstyle\perp}=\{q_{1},q_{2}\}$, $\pi_{3}(t)=q_{3}-A_{3}(t)$, $\pi_{0}=\sqrt{m_{\scriptscriptstyle\perp}^{2}+\pi_{3}^{2}(t)}$, where $m_{\scriptscriptstyle\perp}=\sqrt{1+q_{\scriptscriptstyle\perp}^{2}}$. A typical example is a Sauter pulse,
@@
This equation defines $F$ and $F^{\prime}$ and we have only written out how they depend on the integration variables. We refrain from writing out explicit expressions for the intermediate steps in the calculation, because they are long and not illuminating (the calculations have been done with Mathematica). $F_{1}$ and $F^{\prime}_{1}$ are are nonzero because $\tilde{t}$ is only the saddle point for $t$ and $t^{\prime}$ when the other integration variables are replaced by their saddle-point values.
@@
so all three final-state particles have approximately the same momentum, which is equal to one third of the initial momentum. For the photon-momentum we have a saddle point at $X=0$, so the intermediate photon is approximately on shell. We change variables from ${\bf p}_{1}={\bf p}^{\prime}+\delta{\bf p}_{1}$, ${\bf p}_{2}={\bf p}^{\prime}+\delta{\bf p}_{2}$ and $X=\delta X$ to $\delta{\bf p}_{1}$, $\delta{\bf p}_{2}$ and $\delta X$. All the integration variables with names starting with $\delta$ are $\mathcal{O}(\sqrt{E})$, because $\exp(-h[x_{s}+\delta x]/E)\approx\exp(-[h(x_{s})/E]-[\delta x^{2}/(2E)])$. To simplify we will introduce the average and difference of the electron momenta,
@@
2Re $-2\text{Re }M^{21}\bar{M}^{12}$ (27)
which we call the exchange part. We treat the direct and exchange parts separately. We see from (18) that the exponential and error-function parts of $M^{21}$ is obtained from the corresponding parts of $M^{12}$ by changing sign of $\delta{\bf p}\to-\delta{\bf p}$.
@@ II.1 High-energy/plane-wave limit
where $\tilde{f}^{\prime}=\tilde{f}^{\prime}(\tilde{u})$, $U$ and $V$ are given by (36) and $\tilde{U}$ and $\tilde{V}$ by
@@ II.2 Comparison with plane-wave result
In the following we use $\delta q^{\mu}$ for the momenta as meassured in the rest frame. We still have
@@
For a symmetric, linearly polarized plane wave with a single peak, given by a potential $a_{\mu}=\delta_{\mu,1}\mathcal{A}f(kx)$, where $f(-u)=f(u)$, we have a saddle-point at
@@ II.4 High energy and LCF
We consider arbitrary field shapes, $f(u)$. We can still expand the turning point as in (89), and we again change integration variable from $u=\tilde{u}v$ to $v$. In contrast to the previous section, though, now we can further expand the integrands since $p$ is large. We need at least the next-to-leading order in $1/L$, since the leading order cancels in various places. We normalize the field strength $E$ such that $f^{\prime}(0)=1$, and the frequency scale $\omega$ such that $f^{\prime\prime\prime}(0)=-2$. The reason for choosing $-2$ is because we want the Sauter pulse to be described by $f(u)=\tanh(u)$ rather than, say, $\tanh(u/2^{1/3})$. For the leading and next-to-leading orders ($\mathcal{O}(L^{0})$ and $\mathcal{O}(L^{-2})$) we find
@@
The leading order in $\varphi_{0}$ agrees with the result for trident in a constant-crossed field [5]. We have included the NLO $\mathcal{O}(1/[\chi L^{2}])$ in $\varphi_{0}$ since $\exp(-\text{``something small''}/\chi)$ is not necessarily $\approx 1$. But actually, if one considers a regime where the $1/[\chi L^{2}]$ term is necessary for a good precision, then $\chi\sim 1/L^{2}$ would likely be so small that the overall exponential suppression, $\exp(-16/[3\chi])$, would make the process irrelevant for experiments.
@@
The last term in the square brackets does not look like the corresponding term in (111). This is related to the question of whether the $\gamma\ll 1$ and $p_{0}\gg 1$ limits commute. But in this case we can actually get agreement, although it is somewhat subtle. If the dependence of the spectrum on a single momentum variable $\delta q$ scales as $e^{-\delta q^{2}/\epsilon}$, then a natural interval, e.g. for plotting the spectrum, would be something like $-\sqrt{5\epsilon}\lesssim\delta q\lesssim\sqrt{5\epsilon}$. In the present case we have $\epsilon\propto\chi$ for all momentum variables. However, the scaling of $\epsilon$ with respect to $p$ is not the same for the two round brackets in (65) for $\delta_{1}$ and $\delta p_{3}$. To see how this can be used, we first define
@@
This result can also be seen immediately by noticing that $x$ scales as $p^{0}$, while $y$ scales as $1/p$, so the $x$ term is more exponentially suppressed for generic values of $\delta p_{1}$ and $\delta p_{3}$, which means that the $x$ term basically serves as a delta function for the $y$ term, i.e. one can set $p_{3}\delta p_{1}-p_{1}\delta p_{3}=0$ in $y$. Since (65) has been obtained by expanding to leading order in $1/p$ only, one can replace $y^{2}$ in (65) with $(\delta p_{1}/p_{1}^{\prime})^{2}$, since the difference is higher order in $1/p$ in the natural range of $\delta p_{1}$ and $\delta p_{3}$. But if one only looks at (65) then one might not see any obvious reason for making that replacement. And in (112) one definitely do not see any reason for the replacement, because both $x^{2}/\mathcal{A}^{2}$ and $y^{2}$ scale as $1/L^{2}$ for $\mathcal{A}\sim p\sim L$, so the $x$ term does not serve as a delta function fo the $y$ term. Thus, it is possible to make a replacement in the $p_{0}\gg 1$ limit (65) such that the $\gamma\ll 1$ limit of that “improved” $p_{0}\gg 1$ limit agrees with the correct $p_{0}\sim 1/\gamma\gg 1$ limit (111). However, if we did not already know (111), then there would not have been any obvious reason for making that replacement. And if one does not make that replacement, then one would be led to (112) instead of (111). This raises some questions about how much one can rely on plane-wave results if one is interested in the $p_{0}\sim 1/\gamma\gg 1$ regime.
@@ II.5 Two-step term
where $l_{0}=+|{\bf l}|$ is the on-shell momentum and $\theta(.)$ is the step function. The term with $\theta(x^{0}-y^{0})$ corresponds to a process where the pair-production vertex happens earlier than the Compton vertex, i.e. an electron, a positron and a photon are first produced and then the intermediate photon is later absorbed by the incoming electron. This gives a nonzero contribution, but from semiclassical reasoning we expect it to be subdominant compared to the contribution coming form the $\theta(y^{0}-x^{0})$ term, where the intermediate photon is emitted by the initial electron and later decays into a pair. In fact, by redoing the calculation with
@@
to perform the integral over $\tau=t^{\prime}-t$, we recover exactly the full results in (28) and (29). This shows clearly that the dominant contribution in this semi-classical regime comes from an on-shell intermediate photon which is emitted by the incoming electron and later decays into an electron positron pair. However, the full results still contain coherent contributions, so we want to know if or how these contributions can be neglected, which should be useful for finding more general incoherent-product approximations for more complicated higher-order processes. We will now show that the dominant contribution in the LCF regime can be written as an incoherent product of the probabilities of Compton scattering and Breit-Wheeler pair production. We will follow the ideas in [6] for the construction of a two-step in plane-wave backgrounds. So, we square the amplitude before we perform the $t=x^{0}$ and $t^{\prime}=y^{0}$ integrals. We use $t$ and $t^{\prime}$ for the time integrals coming from the amplitude $M$, and $\bar{t}$ and $\bar{t}^{\prime}$ for the time-integrals coming from the complex conjugate $M^{*}$. Our first step in constructing a two-step term is to drop the $\theta(t-t^{\prime})$ term from (117). The other step function gives on the probability level
@@
We can still perform the time integrals by expanding the exponent around the same saddle point as before. Because of step function (121) we find that, relative to the result without any step functions, we have
@@
where $\tilde{P}_{\rm two}(\epsilon_{\mu}^{({\scriptscriptstyle\parallel})})$ is obtained by only considering an intermediate photon with parallel polarization, and similarly for $\tilde{P}_{\rm two}(\epsilon_{\mu}^{({\scriptscriptstyle\perp})})$. Parallel and perpendicular polarization corresponds ${\bf N}=\{1,0,0,\pm 1\}$. We have

### v2

@@ Abstract
We consider nonlinear trident, $e^{\scriptscriptstyle-}\to e^{\scriptscriptstyle-}e^{\scriptscriptstyle-}e^{\scriptscriptstyle+}$, in various electric background fields. This process has so far been studied for plane-wave backgrounds, using Volkov solutions. Here we first use WKB for trident in time-dependent electric fields, and then for fields which vary slowly in space. Then we show how to use worldline instantons for more general fields which depend on both time and space. For time-dependent fields the WKB approach is at least as simple to use as the worldline approach, but already the relatively modest step of including a slow spatial dependence makes the worldline approach much more efficient.
@@ I Introduction
There continues to be growing progress in strong-field-QED processes, especially those occurring when electrons or photons collide with high-intensity laser fields [1, 2, 3]. One process is the nonlinear trident process, $e^{\scriptscriptstyle-}\to e^{\scriptscriptstyle-}e^{\scriptscriptstyle-}e^{\scriptscriptstyle+}$. Many particles can be produced by a single electron, because factors of $\alpha=1/137$, which would otherwise suppress higher-order processes, are compensated for by large factors due to the high intensity or extended pulse length of the background field. To theoretically model such higher-order processes, one needs to employ various approximations. One approximation, in particular, is to treat higher-order processes as incoherent products of first-order processes. One reason that we are interested in studying trident is that it is the first nontrivial step in the production of more particles while still being sufficiently “simple” so that one can calculate and check the corrections to the incoherent-product approximation. The calculations of these corrections are actually quite nontrivial. In fact, while the simplest part of the correction (the direct part) was calculated already in the 70’s by [4, 5], it was not until [6, 7] that the rest of the correction (the exchange part) was calculated. It turned out that this previously omitted exchange term is in fact on the same order of magnitude as the simpler direct term. Here we should mention that there have also been fully numerical computations of the trident spectrum which did not worry about separating the probability into an incoherent product and corrections to it. A not complete list of papers studying trident is [4, 5, 8, 9, 10, 11, 12, 6, 7, 13, 14, 16, 15, 17, 18].
All these previous studies of trident have in common that they considered the background field to be a plane wave, i.e. with a gauge field given by [^1] We absorb the charge into the definition of the background field, $eF^{\mu\nu}\to F^{\mu\nu}$, and set $c=\hbar=m=1$, where $m$ is the electron mass. $A_{\mu}(kx)$, where $k_{\mu}=\omega(1,0,0,1)$ is the wave vector and $kA=0$. There are good reasons for this. One is the fact that an electron with sufficiently high energy tends to see a quite general background as if it were a plane wave. However, if the electron’s momentum is (approximately) parallel to an electric field, then the electron will not see that field as a plane wave. And if the field is sufficiently strong to significantly bend the electron trajectory, then one can also expect differences from the plane-wave approximation. Then there are of course cases where the electron energy is not high.
@@
In this paper we will use weak-field approximations to calculate the trident probability for non-plane-wave backgrounds. We will start with a WKB treatment of trident in purely time-dependent electric fields, $E(t)=A_{3}^{\prime}(t)$. One might argue about whether this can be seen as going “beyond” plane waves. We are not saying that fields of the form $E(t)$ actually give a more accurate model of realistic laser fields. For one thing, $E(t)$ does not solve Maxwell’s equation, while plane waves do. However, from an analytical point of view, there is a sense in which considering $E(t)$ is going beyond plane waves. The fact that the two field invariants, ${\bf E}^{2}-{\bf B}^{2}$ and ${\bf E}\cdot{\bf B}$, are both zero for plane waves means that the probabilities have less structure. For $E(t)$, on the other hand, one of the field invariants is nonzero, so even before doing any calculations, one can expect the result to have more structure. In fact, results for a general plane wave can be obtained from the results for $E(t)$ by taking the limit where the transverse momentum of the electron is very high. In other words, the plane-wave results are contained in the results for $E(t)$, so in this sense we are going beyond plane waves when studying $E(t)$. Considering $E(t)$ therefore allows us to get a sense of what is neglected when approximating the background as a plane wave, and when it is not possible to do so. We show, in particular, that if we instead take the limit where the momentum of the electron is large but parallel to the electric field, then one does not recover the plane-wave result.
@@
Then we show how to use worldline instantons to study trident in general space-time dependent fields. Both the WKB and the instanton approach calculate the same leading order in a weak-field approximation. In the WKB approach one first has to solve the Dirac equation, i.e. one has to find an approximation of a field, $\psi(x^{\mu})$, which in general depends nontrivially on $t$, $x$, $y$ and $z$ in addition to the field and momentum parameters, and then one has to plug that solution into the Feynman rules and perform the integrals (which one can do with the saddle-point approximation). Note that $\psi(x^{\mu})$ contains more information than what is actually needed for any particular process. In contrast, in the instanton approach one goes directly for the amplitude of the particular process one considers, i.e. without first calculating some intermediate result. Moreover, the instantons are 1D trajectories which permit a more intuitive understanding of these processes. With this in mind, it is perhaps not so surprising that we can calculate things with the instanton approach which, although in principle possible, would be difficult to do with a WKB approach. We will show how to apply the instanton approach for fields which depend on both time and space separately (i.e. not just on e.g. tzt-z as in the plane-wave case). This can be seen as going beyond plane waves both in the analytical sense described above as well as in a more direct sense of more accurately modeling realistic lasers.
@@
  • We want to develop methods to be able to calculate the probability of trident and similar processes for fields which depend on more than one space-time coordinate. Our open-worldline approach seems promising for that.
@@ II Trident in $E(t)$ using WKB
We first consider time-dependent fields of the form $A_{3}(-t)=-A_{3}(t)$, $A_{3}(t)=f(\omega t)/\gamma$, where $f$ is a dimensionless function and $\gamma=\omega/E$ is the Keldysh parameter. In the high-energy limit we will often express things in terms of $a_{0}=1/\gamma$ (the classical nonlinearity parameter) instead of $\gamma$, because $a_{0}$ is more common when dealing with plane-wave backgrounds. We use the following notation for the momentum, $\pi_{\scriptscriptstyle\perp}=q_{\scriptscriptstyle\perp}=\{q_{1},q_{2}\}$, $\pi_{3}(t)=q_{3}-A_{3}(t)$, $\pi_{0}=\sqrt{m_{\scriptscriptstyle\perp}^{2}+\pi_{3}^{2}(t)}$, where $m_{\scriptscriptstyle\perp}=\sqrt{1+q_{\scriptscriptstyle\perp}^{2}}$. A typical example is a Sauter pulse,
@@
This equation defines $F$ and $F^{\prime}$ and we have only written out how they depend on the integration variables. We refrain from writing out explicit expressions for the intermediate steps in the calculation, because they are long and not illuminating (the calculations have been done with Mathematica). $F_{1}$ and $F^{\prime}_{1}$ are nonzero because $\tilde{t}$ is only the saddle point for $t$ and $t^{\prime}$ when the other integration variables are replaced by their saddle-point values.
@@
so all three final-state particles have approximately the same momentum, which is equal to one third of the initial momentum. For the photon-momentum we have a saddle point at $X=0$, so the intermediate photon is approximately on shell [^2] Peaks in the probability associated with intermediate particles which are on or almost on shell have been studied in e.g [20, 21, 22].. We change variables from ${\bf p}_{1}={\bf p}^{\prime}+\delta{\bf p}_{1}$, ${\bf p}_{2}={\bf p}^{\prime}+\delta{\bf p}_{2}$ and $X=\delta X$ to $\delta{\bf p}_{1}$, $\delta{\bf p}_{2}$ and $\delta X$. All the integration variables with names starting with $\delta$ are $\mathcal{O}(\sqrt{E})$, because $\exp(-h[x_{s}+\delta x]/E)\approx\exp(-[h(x_{s})/E]-[\delta x^{2}/(2E)])$. To simplify we will introduce the average and difference of the electron momenta,
@@
$2\text{Re }M^{21}\bar{M}^{12}$ (27)
which we call the exchange part. We treat the direct and exchange parts separately. We see from (18) that the exponential and error-function parts of $M^{21}$ are obtained from the corresponding parts of $M^{12}$ by changing the sign of $\delta{\bf p}\to-\delta{\bf p}$.
@@ II.1 High-energy/plane-wave limit
[…]
@@ II.2 Comparison with plane-wave result
In the following we use $\delta q^{\mu}$ for the momenta as measured in the rest frame. We still have
@@
For a symmetric, linearly polarized plane wave with a single peak, given by a potential $a_{\mu}=\delta_{\mu,1}\mathcal{A}f(kx)$, where f(u)=f(u)f(-u)=f(u), we have a saddle-point at
@@ II.4 High energy and LCF
We consider arbitrary field shapes, $f(u)$. We can still expand the turning point as in (89), and we again change integration variable from $u=\tilde{u}v$ to $v$. In contrast to the previous section, though, now we can further expand the integrands since $p$ is large. We need at least the next-to-leading order in $1/L$, since the leading order cancels in various places. We normalize the field strength $E$ such that $f^{\prime}(0)=1$, and the frequency scale $\omega$ such that $f^{\prime\prime\prime}(0)=-2$. The reason for choosing 2-2 is because we want the Sauter pulse to be described by $f(u)=\tanh(u)$ rather than, say, $\tanh(u/2^{1/3})$. For the leading and next-to-leading orders ($\mathcal{O}(L^{0})$ and $\mathcal{O}(L^{-2})$) we find
@@
The leading order in $\varphi_{0}$ agrees with the result for trident in a constant-crossed field [5]. We have included the NLO $\mathcal{O}(1/[\chi L^{2}])$ in $\varphi_{0}$ since $\exp(-\text{``something small''}/\chi)$ is not necessarily $\approx 1$. But actually, if one considers a regime where the $1/[\chi L^{2}]$ term is necessary for good precision, then $\chi\sim 1/L^{2}$ would likely be so small that the overall exponential suppression, $\exp(-16/[3\chi])$, would make the process irrelevant for experiments.
@@
The last term in the square brackets does not look like the corresponding term in (111). This is related to the question of whether the $\gamma\ll 1$ and $p_{0}\gg 1$ limits commute. But in this case we can actually get agreement, although it is somewhat subtle. If the dependence of the spectrum on a single momentum variable $\delta q$ scales as $e^{-\delta q^{2}/\epsilon}$, then a natural interval, e.g. for plotting the spectrum, would be something like $\sqrt{5\epsilon}\lesssim\delta q\lesssim\sqrt{5\epsilon}$. In the present case we have $\epsilon\propto\chi$ for all momentum variables. However, the scaling of $\epsilon$ with respect to $p$ is not the same for the two round brackets in (65) for $\delta_{1}$ and $\delta p_{3}$. To see how this can be used, we first define
@@
This result can also be seen immediately by noticing that $x$ scales as $p^{0}$, while $y$ scales as $1/p$, so the $x$ term is more exponentially suppressed for generic values of $\delta p_{1}$ and $\delta p_{3}$, which means that the $x$ term basically serves as a delta function for the $y$ term, i.e. one can set $p_{3}\delta p_{1}-p_{1}\delta p_{3}=0$ in $y$. Since (65) has been obtained by expanding to leading order in $1/p$ only, one can replace $y^{2}$ in (65) with $(\delta p_{1}/p_{1}^{\prime})^{2}$, since the difference is higher order in $1/p$ in the natural range of $\delta p_{1}$ and $\delta p_{3}$. But if one only looks at (65) then one might not see any obvious reason for making that replacement. And in (112) one definitely does not see any reason for the replacement, because both $x^{2}/\mathcal{A}^{2}$ and $y^{2}$ scale as $1/L^{2}$ for $\mathcal{A}\sim p\sim L$, so the $x$ term does not serve as a delta function for the $y$ term. Thus, it is possible to make a replacement in the $p_{0}\gg 1$ limit (65) such that the $\gamma\ll 1$ limit of that “improved” $p_{0}\gg 1$ limit agrees with the correct $p_{0}\sim 1/\gamma\gg 1$ limit (111). However, if we did not already know (111), then there would not have been any obvious reason for making that replacement. And if one does not make that replacement, then one would be led to (112) instead of (111). This raises some questions about how much one can rely on plane-wave results if one is interested in the $p_{0}\sim 1/\gamma\gg 1$ regime.
@@ II.5 Two-step term
where $l_{0}=+|{\bf l}|$ is the on-shell momentum and $\theta(.)$ is the step function. The term with $\theta(x^{0}-y^{0})$ corresponds to a process where the pair-production vertex happens earlier than the Compton vertex, i.e. an electron, a positron and a photon are first produced and then the intermediate photon is later absorbed by the incoming electron. This gives a nonzero contribution, but from semiclassical reasoning we expect it to be subdominant compared to the contribution coming from the $\theta(y^{0}-x^{0})$ term, where the intermediate photon is emitted by the initial electron and later decays into a pair. In fact, by redoing the calculation with
@@
to perform the integral over $\tau=t^{\prime}-t$, we recover exactly the full results in (28) and (29). This shows clearly that the dominant contribution in this semi-classical regime comes from an on-shell intermediate photon which is emitted by the incoming electron and later decays into an electron positron pair. However, the full results still contain coherent contributions, so we want to know if or how these contributions can be neglected, which should be useful for finding more general incoherent-product approximations for more complicated higher-order processes. We will now show that the dominant contribution in the LCF regime can be written as an incoherent product of the probabilities of Compton scattering and Breit-Wheeler pair production. We will follow the ideas in [6] for the construction of a two-step in plane-wave backgrounds. So, we square the amplitude before we perform the $t=x^{0}$ and $t^{\prime}=y^{0}$ integrals. We use $t$ and $t^{\prime}$ for the time integrals coming from the amplitude $M$, and $\bar{t}$ and $\bar{t}^{\prime}$ for the time integrals coming from the complex conjugate $M^{*}$. Our first step in constructing a two-step term is to drop the $\theta(t-t^{\prime})$ term from (117). The other step function gives on the probability level
@@
We can still perform the time integrals by expanding the exponent around the same saddle point as before. Because of the step function (121) we find that, relative to the result without any step functions, we have
@@
where $\tilde{P}_{\rm two}(\epsilon_{\mu}^{({\scriptscriptstyle\parallel})})$ is obtained by only considering an intermediate photon with parallel polarization, and similarly for $\tilde{P}_{\rm two}(\epsilon_{\mu}^{({\scriptscriptstyle\perp})})$. Parallel and perpendicular polarization corresponds to ${\bf N}=\{1,0,0,\pm 1\}$. We have